# MuRating: A High Quality Data Selecting Approach to Multilingual Large Language Model Pretraining

### v1

@@ 4.1 Experimental Setups
Dataset construction. We build on the deduplication and heuristic-filtering pipelines of SlimPajama cerebras2023slimpajama and FineWeb lozhkov2024fineweb-edu to assemble a large web-crawl corpus. It comprises 1.5 trillion English tokens plus 3 trillion tokens across 17 additional languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Korean, Portuguese, Russian, Spanish, Thai, Turkish, Vietnamese, Malay, Tagalog), all drawn from Fineweb-2 penedo2024fineweb . We then apply our multilingual rater to assign pairwise-derived quality scores to every document. Although scoring trillions of tokens is compute-intensive, it parallelizes efficiently across GPUs, and batching strategies reduce overhead in practice. Corpus statistics are detailed in Appendix B.1.
@@
Baselines. For English-only experiments, we consider several established data quality raters: QuRater, which selects data based on educational value wettig2024qurating ; AskLLM, which follows the prompt design in sachdeva2024train using Flan-T5-XXL chung2024scaling ; the FineWeb-Edu Classifier [^1] https://huggingface.co/HuggingFaceFW/fineweb-edu-classifier, trained on 450K LLaMA3-70B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct labels to identify educational content; and DCLM [^3] https://huggingface.co/mlfoundations/fasttext-oh-eli5, a fastText classifier trained on high quality dataset to differentiate between informative and low-quality web content.
For the multilingual setting, we adapt the QuRater framework to create QuRater-M, using GPT-4o to annotate multilingual pairs via relative preference, following the same training pipeline as its English counterpart. Additionally, we introduce MuRater(M) as a baseline, which involves sampling multilingual pair data, translating it into English for rating, and projecting the scores back to the original languages. Parallel and cross-lingual pairs are also incorporated to enhance the rater training.
@@
Finally, we include a Uniform baseline for both settings, which randomly samples data (50% more than others). Details are provided in Appendix B.2.
@@
Training Setup. We train a randomly initialized language model based on the LLaMA architecture grattafiori2024llama for a single epoch over the training corpus, with data presented in a randomly shuffled order. The model comprises 1.2 billion parameters and employs a standard transformer architecture vaswani2017attention augmented with rotary position embeddings (RoPE) su2024roformer . To accommodate the multilingual setting, we extend the tokenizer vocabulary through retraining on the multilingual corpus. Comprehensive architectural and tokenizer details are provided in Appendix B.3.
@@
For English-language experiments, we select top scored 200 billion tokens from the full pool of 1.5 trillion tokens across all baseline methods and MuRater. In the multilingual setting, we apply our multilingual rater and the QuRater-M baseline to select the top 10% of tokens per language, resulting in approximately 300 billion tokens in total. These selected multilingual tokens are then combined with the 200 billion English tokens to yield a 500-billion-token training corpus.
@@
Evaluation Benchmarks. We assess the performance of our pretrained models using the lm-evaluation-harness framework eval-harness . For the English-only evaluation, we consider a suite of ten tasks spanning multiple linguistic competencies. These include six reading comprehension benchmarks—ARC-Easy, ARC-Challenge clark2018think , SciQ welbl-etal-2017-crowdsourcing , LogiQA liu2020logiqa , TriviaAQjoshi2017triviaqa and BoolQ clark2019boolq ; four commonsense reasoning tasks—HellaSwag zellers2019hellaswag , PIQA bisk2019piqa , OpenBookQA OpenBookQA2018 and WinoGrande sakaguchi2019winogrande ; and two knowledge-intensive tasks—Natural Questions (NQ) kwiatkowski2019natural and MMLU hendrycks2021measuring . For MMLU, we follow alzahrani2024benchmarks and employ the lighteval variant to ensure more consistent and reliable comparisons.
For the multilingual evaluation, we utilize translated versions of several English benchmarks in addition to multilingual-native datasets han2025mubenchassessmentmultilingualcapabilities . The evaluation suite includes translated ARC-Easy, ARC-Challenge, and HellaSwag, MMLU, StoryCloze mostafazadeh2016corpus along with XCOPA ponti2020xcopa , XNLI conneau2018xnli , XWinograd tikhonov2021s , BMLAMA qi-etal-2023-cross and FLORES goyal2022flores , We further incorporate native language benchmarks—CMMLU li2024cmmlu , VMLU  [^4] https://vmlu.ai/, IndoMMLU koto-etal-2023-indommlu , JMMLU [^5] https://huggingface.co/datasets/nlp-waseda/JMMLU, and AMMLU [^6] https://huggingface.co/datasets/Hennara/ammlu—to construct a localized multilingual variant of the MMLU test set, denoted as MMLU_L. These datasets collectively evaluate the model’s capabilities in cross-lingual comprehension, reasoning, and translation resilience. A detailed overview of the benchmarks and their language coverage is provided in Appendix B.4.
@@ 4.2.1 Multilingual Results
Table 1: Performance of different data selection strategies across downstream tasks under mixing 200B English and 300B multilingual tokens. MuRater(M) denotes training with multilingual pairs translated into English for scoring, while MuRater(E) uses rated English data translated into multilingual pair form. Best results within each setting are shown in bold.
Selection Method MMLU_L ARC_C_ML ARC_E_ML Flores Hellaswag_ML MMLU_T XCOPA XNLI StoryCloze_ML XWinograd BMLAMA Average
Uniform(+50% data) 29.98 28.74 49.03 46.66 44.56 27.83 64.60 42.33 69.28 76.40 48.55 48.00
QuRater-M 30.86 33.89 57.53 46.60 46.77 29.18 62.97 44.39 65.86 71.21 45.84 48.65
MuRater(M) 31.86 34.26 58.21 47.43 46.54 29.28 64.43 42.07 67.08 72.92 50.13 49.47
MuRater(E) 31.96 35.01 58.98 47.27 47.11 29.40 65.73 44.46 67.67 74.13 52.97 50.43
The results in Table 1 and Figure 2 demonstrate that MuRater consistently outperforms the multilingual baseline QuRater-M across most benchmarks by 1.8 percent, particularly excelling on knowledge-intensive and reading comprehension tasks such as ARC and MMLU. This indicates that MuRater is more effective in selecting high-quality multilingual pretraining data. Among its two variants, MuRater(E), which translates rated English data into multilingual counterparts, achieves better performance than MuRater(M), suggesting that transferring rating capabilities from English to other languages via translation is a more efficient strategy. Despite its overall advantage, MuRater shows relatively weaker results on narrative and context-rich tasks like StoryCloze-ML and XWinograd, where the Uniform baseline performs better. These tasks require nuanced contextual comprehension and familiarity with diverse narrative styles, which are better captured by larger and more varied training data. MuRater’s preference for fact-based, educationally relevant content boosts performance in QA-style evaluations but may lead to underrepresentation of narrative-rich samples. Furthermore, the Uniform baseline benefits from 50% more data, which contributes to its relative strength on narrative-oriented tasks. Detailed language-specified performance breakdowns are provided in Appendix D.
@@ 4.2.2 English-only Results
Table 2: Performance of different selection method over all different downstream tasks. Best results of each task category is marked in black. Detailed results are performed in Appendix D.
@@
The results in Table 2 indicate that our proposed rater successfully consolidates the strengths of existing rating methodologies, leading to consistent improvements in pretrained model performance across all categories of evaluation tasks. Baseline comparisons reveal that each selection method exhibits distinct preferences for data, which translate into varying levels of effectiveness on different downstream tasks. For instance, as shown in Figure 3, DCLM yields strong results on HellaSwag but underperforms on ARC-Challenge. Conversely, QuRater achieves competitive performance on ARC-Challenge but demonstrates poor results on TriviaQA. In contrast, our rater integrates the advantages of these methods and achieves robust performance across nearly all benchmarks, outperforming other data selection baselines by margins ranging from 1 to 3.4 percent. The model trained with our rater consistently achieves superior results on all tasks throughout the training process. This indicates more stable and efficient learning, further validating the effectiveness of our data selection approach in enhancing the quality of LLM pretraining.
(a) Average performance
(b) ARC-Challenge
(c) HellaSwag
(d) TriviaQA
@@ 4.3.1 Effectiveness of Cross-Lingual and Parallel Pair Integration
Incorporating cross-lingual pairs and parallel translations during training significantly improves the consistency of quality scoring across languages. To validate this, we assess multilingual raters on parallel corpora—semantically equivalent texts in different languages. In an ideal scenario, a language-agnostic rater would assign same scores for parallel language texts, yielding a slope of one and minimal mean squared error (MSE) between the score sequences of the corresponding texts. As shown in Figure 4, our alignment-based training yields models with lower MSE and slopes closer to one, demonstrating improved cross-lingual consistency.
These findings highlight the importance of modeling interlingual relationships. By leveraging cross-language comparisons and parallel data, the rater learns language-invariant quality standards, enabling more reliable multilingual evaluation. A qualitative case study in Appendix E further supports this, showing that high-rated texts consistently exhibit greater fluency, coherence, and instructional value across languages. We further examine how the selected data is distributed across semantic domains in different languages. Detailed results are provided in Appendix B.1.
In addition, we sample 10,000 annotated samples of nine languages using both aligned and unaligned raters. We compute the Kendall Tau coefficients between the resulting score sequences and visualize them in Figure 4(d). The strong correlation indicates that alignment mainly standardizes absolute score values across languages without affecting the relative ranking. Moreover, selecting the top 10% high-quality samples from both configurations yields a 92% overlap, confirming that alignment preserves core evaluative behavior while enabling more consistent cross-lingual interpretation.
@@
(d) Kendall Tau Corr.
Figure 4: Scatter plots of scores assigned by multilingual raters to 10,000 parallel documents across various languages. Green points represent ratings from raters trained with alignment using parallel and cross-lingual pairs, while blue points indicate scores from unaligned raters. The Kendall Tau matrix illustrates the rank correlation between the two raters’ scoring outputs of 10,000 documents of different languages.
4.3.2 Translation Quality Asseement
Figure 5: Translation average scores of various languages.
We assess the quality of our translations through human evaluation. Expert annotators are provided with 50 pairs of source and translated texts and asked to rate translation quality on a scale from 1 to 5: scores 1 indicate severely flawed translations, 2 corresponds to incorrect translation with notable errors, 3 denotes generally correct translations with some errors, 4 represents good translations with minor issues, and 5 corresponds to near-perfect translations. The evaluation criteria are detailed in Appendix C.1. As shown in Figure 5, the overall translation quality of GPT-4o is high, with most languages achieving average scores above 4. Notably, performance on Japanese and Thai is comparatively lower, though still above 3.5, suggesting acceptable translation quality for these languages.
@@ 4.3.3 Comparison Between Pairwise and Pointwise Score Transfer
We examine the relative effectiveness of pairwise versus pointwise judgment methods for transferring English scoring capabilities to multilingual settings. Based on the translation quality evaluations, we select two high-performing languages, Arabic and Spanish, for the study. Specifically, we translate 200 English text pairs into Arabic and 200 pairs into Spanish. Each dataset is then annotated by GPT-4o using both pairwise and pointwise scoring strategies. For pointwise annotation, GPT-4o assigns quality scores on a 1–10 scale. The scoring prompts of both methods explicitly instruct GPT-4o to evaluate based on content quality alone, irrespective of language and are detailed in Appendix C.3. Each text or pair is scored 20 times, and the average is used as the final score. Given identical content across different languages, the ideal scenario is that a consistent model and prompt should yield nearly identical scores, regardless of the language and score strategies.
As shown in Figure 6, pointwise scores exhibit greater variability across languages, particularly in the mid-quality range (scores between 3 and 6), despite relative stable assessments at the high and low ends. In contrast, pairwise judgments demonstrate high consistency across languages, with only minor pair-score discrepancies observed. These findings suggest that while translation quality is generally sufficient, minor translation biases can still influence pointwise ratings. The pairwise approach, however, proves more robust to such variation, supporting its suitability for reliably transferring English scoring capabilities to multilingual contexts.
@@ 5 Conclusion
We have introduced MuRating, a unified and scalable approach to multilingual data selection that leverages high-quality English quality signals and transfers them across seventeen target languages. By first aggregating multiple English raters through a pairwise Bradley–Terry framework and then projecting those judgments via translation into monolingual, cross-lingual, and parallel text pairs, MuRating yields a single language-agnostic scorer capable of annotating trillions of tokens with minimal compute. When used to curate a 500 billion-token pretraining corpus (200 billion English + 300 billion multilingual), MuRating outperforms strong baselines—including uniform sampling (+50 % data), QuRater, AskLLM, and DCLM—on a suite of twelve English benchmarks and a diverse multilingual evaluation. Our experiments further demonstrate that pairwise translation supervision is more robust than pointwise scoring and cross-lingual and parallel-pair regularization greatly improve score consistency. While MuRating particularly boosts performance on knowledge-intensive and QA-style tasks, it underselects narrative-rich content, pointing to future avenues in genre-aware filtering and dynamic sampling schedules. We will release our prompts, code, and scored datasets to facilitate broader research into multilingual data quality and more equitable LLM development.
@@
Currently we’ve tried it on 17 tongues—there’s plenty more to explore. Relying on GPT-4o brings its own quirks (and blind spots), and our taste for factual content means we don’t give stories and creative writing their fair shake. We’ve only tested on a 1.2 B-parameter model, and scoring trillions of tokens still eats up a ton of GPU hours. In the future, we’ll lean on better translations, expand to more languages, and add smarter sampling to capture a richer, more diverse mix of text.
Appendix A Details of MuRater Model
A.1 Different Annotation Method
GPT annotation We adopt the educational value prompt criteria from QuRating [55] as our annotation prompt for GPT-4o-08-06, as detailed below. This prompt is used to annotate a total of 300,000 document pairs. For each pair, we randomly extract a segment of $n$ tokens—based on the LLaMA tokenizer [50]—where $n$ is sampled from a uniform distribution $n\sim\text{Uniform}[256,512]$ in 50% of cases, and fixed at 512 tokens otherwise. Annotation involves generating 20 predictions of either “A” or “B” per criterion and document pair (in either order). The total cost of dataset creation amounts to \$9,740.

### v2

@@ 4.1 Experimental Setups
Dataset construction. We build on the deduplication and heuristic‐filtering pipelines of FineWeb-2 penedo2025fineweb2pipelinescale to assemble a large web‐crawl corpus. It comprises 1.5 trillion English tokens plus 3 trillion tokens across 17 additional languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Korean, Portuguese, Russian, Spanish, Thai, Turkish, Vietnamese, Malay, Tagalog). We then apply MuRater and other baselines to assign quality scores to every document. Although scoring trillions of tokens is compute-intensive, it parallelizes efficiently across GPUs, and batching strategies reduce overhead in practice. Corpus statistics are detailed in Appendix 7.1.
@@
Baselines. For the English experiments, we train and evaluate the model using only English pairwise data, comparing it against several established data-quality raters: QuRater, which selects data based on educational value wettig2024qurating; AskLLM, which follows the prompt design in sachdeva2024train using Flan-T5-XXL chung2024scaling; the FineWeb-Edu Class ifier [^2] https://huggingface.co/HuggingFaceFW/fineweb-edu-classifier, trained on 450K LLaMA3-70B-Instruct  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct labels to identify educational content; and DCLM [^4] https://huggingface.co/mlfoundations/fasttext-oh-eli5, a fastText classifier trained on high quality dataset to differentiate between informative and low-quality web content.
For the multilingual experiments, we extend the QuRater framework wettig2024qurating to build QuRater-M, which employs GPT-4o to annotate multilingual pairs via relative preference, following the same training pipeline as its English counterpart. We further include datasets from HPTL de2024new and FineWeb-2 penedo2025fineweb2pipelinescale for comprehensive comparison, and evaluate against FineWeb2-HQ messmer2025enhancing. In addition, we analyze the difference between MuRater(E) and MuRater(M): MuRater(M) scores multilingual pairs that have been translated into English, whereas MuRater(E) starts from rated English data and translates it into multilingual pair and cross-lingual pair form. Most multilingual evaluations cover 17 languages plus English, while the experiment involving FineWeb2-HQ is restricted to 13 overlapping languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Portuguese, Russian, Spanish, Turkish, and Vietnamese) to ensure fair comparison.
@@
Finally, we include a Uniform baseline for both settings, which randomly samples 50% more data than the other methods, following the setup of QuRating wettig2024qurating. Details are provided in Appendix 7.2.
@@
Training Setup. We train a randomly initialized language model based on the LLaMA architecture grattafiori2024llama for a single epoch over the training corpus, with data presented in a randomly shuffled order. For most experiments, the model comprises 1.2 billion parameters and employs a standard transformer architecture vaswani2017attention augmented with rotary position embeddings (RoPE) su2024roformer. To accommodate the multilingual setting, we extend the tokenizer vocabulary through retraining on the multilingual corpus. Comprehensive architectural and tokenizer details are provided in Appendix 7.3.
@@
Building on this setup, we construct the training corpora for both English and multilingual experiments. For the English setting, we select the top-scored 200 billion tokens from the full pool of 1.5 trillion tokens across all baseline methods and MuRater. In the multilingual setup, we apply all methods to score and select the top 10% of tokens within each language, yielding roughly 300 billion tokens in total. These multilingual tokens are then combined with the 200 billion English tokens to form a unified 500-billion-token pretraining corpus.
To further assess robustness and scalability under varied training conditions, we additionally conduct experiments with a 7B-parameter model sharing the same LLaMA architecture as the 1.2B model. This larger model is pretrained on 1T tokens, with 16.5% multilingual data selected by either MuRater or QuRater-M, while the remaining 83.5%—comprising English, code, and math data—remains identical across both setups. The multilingual portion follows the same language distribution as in the main experiments, enabling a more comprehensive evaluation of generalization across heterogeneous data sources.
@@
Evaluation Benchmarks. We assess the performance of our pretrained models using the lm-evaluation-harness framework eval-harness. For the English-only evaluation, we consider a suite of ten tasks spanning multiple linguistic competencies. These include six reading comprehension benchmarks—ARC-Easy, ARC-Challenge clark2018think, SciQ welbl-etal-2017-crowdsourcing, LogiQA liu2020logiqa, TriviaAQjoshi2017triviaqa and BoolQ clark2019boolq; four commonsense reasoning tasks—HellaSwag zellers2019hellaswag, PIQA bisk2019piqa, OpenBookQA OpenBookQA2018 and WinoGrande sakaguchi2019winogrande; and two World Knowledge tasks—Natural Questions (NQ) kwiatkowski2019natural and MMLU hendrycks2021measuring. For MMLU, we follow alzahrani2024benchmarks and employ the lighteval variant to ensure more consistent and reliable comparisons.
For the multilingual evaluation, we utilize translated versions of several English benchmarks han2025mubenchassessmentmultilingualcapabilities alongside multilingual-native datasets. Similar to English benchmarks, we divide the task into three categories, including reading comprehension, commonsense reasoning, and world knowledge understanding. Reading comprehension is evaluated using translated versions of ARC-Easy and ARC-Challenge clark2018think, StoryCloze mostafazadeh2016corpus, and XNLI conneau2018xnli, which assess contextual understanding and inference. Commonsense reasoning is tested with HellaSwag zellers2019hellaswag, XCOPA ponti2020xcopa, and XWinograd tikhonov2021s, focusing on event causality, semantic plausibility, and everyday reasoning across languages. World Knowledge evaluation includes MMLU hendrycks2021measuring, BMLAMA qi-etal-2023-cross, and FLORES goyal2022flores, which examine factual knowledge, translation quality, and multilingual alignment.
To strengthen the evaluation of language-specific knowledge, we also incorporate localized MMLU variants—CMMLU li2024cmmlu, VMLU [^5] https://vmlu.ai/, IndoMMLU koto-etal-2023-indommlu, JMMLU [^6] https://huggingface.co/datasets/nlp-waseda/JMMLU, and AMMLU [^7] https://huggingface.co/datasets/Hennara/ammlu—to construct a region-specific multilingual extension of MMLU, denoted as MMLU_L and add to the world knowledge category. Together, these benchmarks provide a comprehensive evaluation of cross-lingual comprehension, commonsense inference, and knowledge-grounded reasoning, enabling a holistic assessment of multilingual LLM performance. Detailed information of benchmark statistics and language coverage is provided in Appendix 8.
@@ 4.2.1 Multilingual Results
Table 1: Results on multilingual benchmarks with different training setups. Best results within each setting are shown in bold
Selection Method Reading Comprehension (5 tasks) Commonsense Reasoning (2 tasks) World Knowledge (4 tasks) Average (11 tasks)
18 Languages Results
Uniform 53.16 54.58 38.25 48.66
HPLT-2 50.38 49.77 36.96 45.70
FineWeb-2 50.83 52.48 35.53 46.28
QuRater-M 54.58 54.87 38.12 49.19
MuRater(M) 54.91 55.48 39.68 50.02
MuRater(E) 56.05 56.42 40.40 50.96
13 Languages Results
FineWeb2-HQ 53.05 55.54 38.31 48.97
MuRater(E) 55.95 58.30 41.17 51.81
7B Model Results
QuRater-M 61.96 63.28 43.31 56.18
MuRater 62.78 64.40 44.50 57.23
As shown in Table 1 and Figure 2, MuRater substantially outperforms existing multilingual baselines across nearly all evaluation categories and settings. Under the 18-language configuration, MuRater(E) achieves the highest category-averaged scores in all three categories, outperforming all baselines. These consistent gains highlight MuRater’s capacity to identify high-quality, semantically rich, and educationally valuable text, even when faced with heterogeneous multilingual inputs. In particular, the large improvements in reasoning-oriented benchmarks (e.g., ARC and MMLU) suggest that MuRater selects examples with deeper conceptual structure and higher linguistic clarity, thereby enhancing model comprehension and reasoning generalization.
The performance gap between MuRater(M) and MuRater(E) further highlights the advantages of English-anchored training. MuRater(E) leverages English-rated pairs and projects the resulting preferences into the multilingual space, providing more stable and transferable supervision signals. We attribute this benefit to the broader topical and stylistic diversity of English corpora, which expose the rater to richer linguistic variation and more representative pairwise patterns during training. As a result, MuRater(E) learns to generalize beyond language corpus boundaries and more effectively capture shared semantic dimensions across languages, yielding stronger cross-lingual alignment and greater robustness to translation-induced noise.
On the 13-language subset, MuRater(E) continues to outperform FineWeb2-HQ by roughly 3 points on average, achieving leading results across all three evaluation dimensions. When scaled to the 7B model trained on 1T tokens, MuRater maintains its superiority, reaching higher performance in all three respective categories—demonstrating consistent gains across both model sizes and data distributions. These results confirm that MuRater generalizes effectively to diverse linguistic environments and evaluation conditions, offering a more scalable and reliable framework for multilingual data quality estimation. Detailed results on each language are shown in Appendix 9
@@ 4.2.2 English-only Results
Table 2: Performance of different selection method over all different downstream tasks. Best results of each task category is marked in black. Detailed results are performed in Appendix 9.
@@
The results in Table 2 indicate that our proposed rater successfully consolidates the strengths of existing rating methodologies, leading to consistent improvements in pretrained model performance across all categories of evaluation tasks. Baseline comparisons reveal that each selection method exhibits distinct preferences for data, which translate into varying levels of effectiveness on different downstream tasks. For instance, as shown in Figure 3, DCLM yields strong results on HellaSwag but underperforms on ARC-Challenge. Conversely, QuRater achieves competitive performance on ARC-Challenge but demonstrates poor results on TriviaQA. In contrast, our MuRater integrates the advantages of these methods and achieves robust performance across nearly all benchmarks, outperforming other data selection baselines by margins ranging from 1 to 3.4 percent. The model trained with our rater consistently achieves superior results on all tasks throughout the training process. This indicates more stable and efficient learning, further validating the effectiveness of our data selection approach in enhancing the quality of LLM pretraining.
(a) ARC-Challenge
(b) HellaSwag
(c) TriviaQA
@@ 4.3.1 Effectiveness of Cross-Lingual and Parallel Pair Integration
Incorporating cross-lingual pairs and parallel translations during training significantly improves the consistency of quality scoring across languages. To validate this, we assess multilingual raters on parallel corpora—semantically equivalent texts in different languages. As shown in Figure 4, our alignment-based training produces models with lower mean squared error (MSE) and slopes closer to one, indicating stronger cross-lingual consistency. In an ideal case, a language-agnostic rater would assign identical scores to parallel texts across languages, resulting in a slope of one and minimal MSE between their score sequences.
These findings highlight the importance of modeling interlingual relationships. By leveraging cross-language comparisons and parallel data, the rater learns language-invariant quality standards, enabling more reliable multilingual evaluation. A qualitative case study in Appendix 10 further supports this, showing that high-rated texts consistently exhibit greater fluency, coherence, and instructional value across languages. We further examine how the selected data is distributed across semantic domains in different languages. Detailed results are provided in Appendix 7.1.
@@
Figure 4: Scatter plots of scores assigned by multilingual raters to 10,000 parallel documents across various languages. Green points represent ratings from raters trained with alignment using parallel and cross-lingual pairs, while blue points indicate scores from unaligned raters.
@@ 4.3.2 Comparison Between Pairwise and Pointwise Score Transfer
We examine the relative effectiveness of pairwise versus pointwise judgment methods for transferring English scoring capabilities to multilingual settings. Based on the translation quality evaluations, we select two high-performing languages, Arabic and Spanish, for the study. Specifically, we translate 200 English text pairs into Arabic and 200 pairs into Spanish. Each dataset is then annotated by GPT-4o using both pairwise and pointwise scoring strategies. For pointwise annotation, GPT-4o assigns quality scores on a 1–10 scale. The scoring prompts of both methods explicitly instruct GPT-4o to evaluate based on content quality alone, irrespective of language and are detailed in Appendix 6.6. Each text or pair is scored 20 times, and the average is used as the final score. Given identical content across different languages, the ideal scenario is that a consistent model and prompt should yield nearly identical scores, regardless of the language and score strategies.
As shown in Figure 5, pointwise scores exhibit considerable variability across languages, particularly in the mid-quality range (scores between 3 and 6), despite relatively stable assessments at the high and low ends. Ideally, all points should align closely with the $y=x$ (slope = 1) line, which would indicate identical score for semantically equivalent content across languages. In contrast, pairwise judgments display strong cross-lingual consistency, with only minor deviations from this ideal alignment. These observations suggest that while translation quality is generally adequate, subtle translation biases can still influence absolute (pointwise) ratings. The pairwise approach, however, demonstrates greater robustness to such variation, underscoring its effectiveness for reliably transferring English scoring behavior to multilingual contexts.
@@ 5 Conclusion
We introduced MuRating, a scalable multilingual data selection framework that aggregates multiple English raters via a Bradley–Terry pairwise model and transfers these judgments through translation to train a single multilingual MuRater over monolingual, cross-lingual, and parallel pairs. Applied to large web corpora, MuRater is used to pretrain both 1.2B- and 7B-parameter LLaMA-architecture models and delivers consistent improvements over strong baselines (QuRater, FineWeb2-HQ, AskLLM, DCLM) on multiple English and multilingual benchmarks. Ablation results further demonstrate the effectiveness of incorporating cross-lingual and parallel pairs, and confirm that pairwise supervision provides more stable multilingual scoring than pointwise methods. Analyses across translation fidelity and data composition, together with results at two model scales, indicate that MuRating is effective and scalable for large-scale multilingual data curation and yields reliable performance gains across evaluation settings.
@@
Our current study focuses on 17 target languages excluding English, leaving substantial room for broader linguistic coverage. The reliance on GPT-4o introduces potential biases and idiosyncrasies inherent to proprietary large language models. Moreover, since the English raters used in our framework primarily focused on factual and informational content, our auto-rater exhibits limited performance on narrative and creative domains. While the proposed approach performs well on language-specific benchmarks, further research could explore language-specific rater designs or culturally aligned data selection strategies to better capture the unique characteristics of each language. Future work will also aim to incorporate higher-quality translations, expand to a wider range of languages, and develop adaptive data sampling techniques to enrich the diversity and representativeness of the curated multilingual corpus.
6 Details of MuRater Model
6.1 Different Annotation Method
GPT annotation We adopt the educational value prompt criteria from QuRating [wettig2024qurating] as our annotation prompt for GPT-4o-08-06, as detailed below. This prompt is used to annotate a total of 300,000 document pairs. For each pair, we randomly extract a segment of $n$ tokens—based on the LLaMA tokenizer [touvron2023llama]—where $n$ is sampled from a uniform distribution $n\sim\text{Uniform}[256,512]$ in 50% of cases, and fixed at 512 tokens otherwise. Annotation involves generating 20 predictions of either “A” or “B” per criterion and document pair (in either order). The total cost of dataset creation amounts to \$9,740.